OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (anonymous):

set y=0 to find x intercept it will yield x=-3/2 set x=0 to find y intercept, it will yield y=-3

OpenStudy (anonymous):

@psi9epsilon, How do I do that?

OpenStudy (anonymous):

well we need to find intercepts, so, to begin with say we want to find x intercept, then in the function put y=0 that should give 0=-2x-3, which works out to x=-3/2 similarly do it for y intercept